# Technologies for Modulation of Visible and Near-UV Light and their Applications

### v1

@@ Abstract
Control over the amplitude, phase, and spatial distribution of visible-spectrum light underlies many technologies, but commercial solutions remain bulky and require high control power. Recent materials and fabrication advances have spurred interest in visible-light integrated photonics. In this review, we discuss three application spaces which rely on control of visible light - quantum control and measurement of neutral atom and trapped-ion atomic systems, augmented-reality displays, and measurement and control of biological systems. We then review the commercial dynamic surfaces and bulk systems which currently provide visible-light modulation before discussing the current state-of-the-art integrated solutions. Throughout the review we focus on speed, control power, size, optical bandwidth, and technological maturity when comparing technologies.
@@
Control over light has revolutionized fields from communication to the semiconductor industry. In the 19th century, Pockels and others investigated how external electric and strain fields could change the optical properties of materials, enabling a host of bulk optical elements for manipulating light. These bulk modulators are the current workhorse for light modulation with near-unity efficiencies and modulation speeds up to GHz. However, these components provide limited spatial control, are generally large, and require high drive power per channel.
@@
Since the early 1990s, miniaturized and integrated systems have been recognized as the future of active light control [1] and integrated photonics and controllable surfaces promise to reduce both the size and power requirements of visible-light modulation devices. Indeed, for light with frequencies below the bandgap of Silicon (Si) (1.12 eV, $\lambda>1100$ nm) great progress has been made in developing compact, fast modulators based on integrated photonics [2, 3, 4]. By leveraging the trillion-dollar CMOS industry, Si integrated photonics has been able to move from proof-of-principle demonstrations to deployment in data centers and high-rate intercontinental communication [5]. However, the miniaturization and commercialization of devices to control visible light has so far lagged behind.
@@
Control of visible light at low per-channel drive power, compact form-factor, and high frequencies will enable the growth of a wide variety of technologies. In particular, control and measurement of quantum systems, augmented-reality displays, and imaging and control of biological systems are three application areas that have enjoyed significant growth in the last decade, but which are hampered by the capabilities of current commercially-available visible-light modulation. This explosion of application areas coupled with materials and processing developments has accelerated the research in and development of technologies for visible-light modulation. Here, we review the needs of these three application spaces and the current state-of-the-art for visible-light modulation technologies. We aim to provide a review of the optical state-of-the-art for researchers in application fields who are interested in scaling up modulation of visible light and a review of the application space to motivate and focus further research in scalable visible light modulation platforms.
@@
In Sec. I, we discuss the needs of three application spaces in detail - quantum control and readout of atomic qubits, augmented-reality displays, and control of biological systems. These technologies serve as an anchor point for the figures of merit of the various visible-light modulation platforms discussed in the following sections. In Sec. II, we discuss reflective and transmissive controllable surfaces. These, mainly commercial, solutions enable current application systems but suffer from high per-channel power and size, and often cannot reach the necessary modulation speeds. In Sec. III, we discuss integrated solutions to visible-light modulation. These nascent technologies need significant progress to reach the maturity for systems-level integration, but hold great promise for the discussed applications. Finally, we conclude with a perspective on future research directions, including the peripheral infrastructure needed to transition these technologies out of the laboratory and into the field.
@@ I Applications
While we focus our discussion to the above applications, there are a plethora of other potential applications which may be enabled by improved modulation of visible wavelengths. For instance, both classical and quantum optical computing would benefit from moving to visible wavelengths to increase detection efficiency and reduce component size. Similarly, most LIDAR demonstrations use telecommunications light, but visible LIDAR can outperform infrared systems in underwater applications due to reduced scattering [6, 7, 8].
@@ I.1 Quantum Technologies
Quantum technologies promise improvements over their classical counterparts in communication [9], sensing [10], timekeeping [11, 12], and computation [13]. However, to provide utility beyond proof-of-principle demonstrations it will be necessary to reduce the size, weight, and power of all components while simultaneously improving stability and capability. Neutral and charged (ionic) qubits are two of the leading platforms for quantum information processing (QIP) and both rely on ultraviolet, visible, and near infrared light for control and measurement. Although we do not discuss them explicitly here, a broad range of solid state atomic-like defects, for example in diamond [14, 15] or SiC [16], also rely on visible light modulation, with requirements similar to neutral atoms and trapped ions.
The majority of experimental demonstrations of QIP have relied on bulk crystals to modulate the light needed to control and readout the qubit state. However, to scale these systems to the millions of qubits necessary for fault-tolerant quantum computation [17], it will be necessary to actively modulate the amplitude, frequency, phase, and often position of millions of channels simultaneously at wavelengths across the UV and visible spectrum. This is untenable with current solutions. Similarly, deployable quantum sensors, clocks, and quantum network nodes will need compact, stable optical control. It is thus clear that integrated photonics will play a critical role in future quantum technologies [18, 19, 20].
@@ I.1.1 Trapped Ions
Figure 1: Electronic level structure for ${}^{40}$${}^{+}$, ${}^{137}$${}^{+}$, and ${}^{171}$${}^{+}$ as three representative ions, and for ${}^{87}$b and ${}^{171}$b as two representative neutral-atoms.
Trapped ions are a promising platform for quantum simulation [21, 22], quantum sensing [23, 24, 25], optical clocks [26], and gate-based quantum computation [27, 28, 29]. Indeed, the state-of-the-art systems provide the highest reported quantum volume to date [30, 31]. In trapped-ion quantum technologies, atomic ions (generally singly charged) are trapped in ultrahigh vacuum using electromagnetic fields. Quantum information is stored in the coherence between two stable electronic states of each ion. Each ion qubit is by nature identical and can be well-isolated from the environment as they are trapped far from surfaces. Furthermore, ions trapped together in the same confining potential share motional modes that can be used to directly mediate entanglement between qubits [32], leading to natural all-to-all connectivity. Entanglement between ions in different potentials can be mediated by either physically shuttling the ions between trapping regions [31] or optically [33].
Fig. 1 shows relevant transitions for controlling ${}^{40}$${}^{+}$ and ${}^{137}$${}^{+}$, and ${}^{171}$${}^{+}$, three commonly used ions. Other ions and isotopes exhibit similar level structures, though with different transition frequencies. As evidenced in Fig. 1, control of even a single species can span the spectrum from the UV to the near infrared. Additionally, it is likely that a final trapped-ion QIP architecture will employ multiple ion species [31] or encodings [34] for mid-circuit cooling and measurement.
[…]
@@
Recently, piezo-actuated waveguide modulators (see Sec. III.3) were used to control the state of a ${}^{40}$${}^{+}$ trapped-ion qubit with MHz switching speeds and nearly 40 dB on-off extinction [58]. With the integrated modulators they measured 99.7 % single qubit gate fidelity with randomized benchmarking and performance that nearly matched the performance when using standard acoustic optic modulation. However, they found that the integrated modulators were significantly less stable, highlighting that significant device and systems-level improvements are needed to reach the stringent requirements needed for quantum control.
I.1.2 Neutral-Atoms in Optical Tweezer Array
@@
Neutral-atom arrays are an another leading platform for quantum metrology [59, 60, 61, 62], simulations [63, 64, 65, 66, 67, 68, 64], information processing [69, 70, 71, 72, 73]. Recently, it has also enabled fault-tolerant quantum processing based on tens of error-corrected logical qubits [73]. As laser-cooled and then loaded into programmable arrays of optical tweezers [74], the size of neutral-atom arrays scales up to defect-free hundreds to thousands of atoms [66, 75, 76, 77, 78]. Since the atoms are identical and well-isolated from environment, they provide outstanding memory qubits for quantum technologies.
@@
In neutral atom QIP, lasers are used to trap and control the position of individual neutral-atom qubits and mediate single and multi-qubit gates. For entangling gates, laser beams are used to selectively excite target atoms to highly excited Rydberg states, eanbling mutual interactions via the Rydberg blockade effect [79, 80]. Nearby-atoms cannot be simultaneously promoted to the Rydberg state, implementing the conditional-NOT [81, 70, 69, 80] and Toffoli [82] gates between the memory qubits. This effect can also be used to control the system’s many-body Hamiltonian for quantum simulation and optimization [83, 65, 66, 71]. Moreover, the dynamic programming of tweezer positions allows for transporting the entangled qubits across the arrays, featuring the reconfigurable, all-to-all connectivities essential for efficiently implementing hard problems [71, 84], quantum error-correction codes [69, 85] and fault-tolerant information processing [86, 87, 73].
@@
Controlling neutral-atom qubits uses multiple colors of laser light (Fig. 1). For instance, an array of green laser foci creates dipole trap arrays to which ${}^{171}$b qubits are loaded [88]. Driving the ground-to-Rydberg transition of the Rubidium uses 297 nm (the combination of 480 and 780 nm or 420 and 1013 nm) lasers for the single- (two-) photon transition [89, 90, 83]. Laser cooling and trapping of neutral atoms requires kHz timing control, and the timescales for quantum gates typically ranges from 1 MHz [63, 68] to 100 MHz [80, 90, 91] depending on the application. Moreover, since the atom arrangement determines the qubit connectivity [69, 71, 84] the dynamic control of spatial modes combined with the fast temporal modulation (e.g., controlling the position of the lasers providing the trapping potentials) is essential for assembling atom arrays and implementing local qubit control capabilities [73].
@@
Implementing the required control capabilities utilizes different modulation mechanisms. For optical tweezers, liquid crystal on silicon (LCoS) spatial light modulators (Section III.4), mechanical mirror arrays (Section II.1), and acousto-optic modulators (Section II.2.1) are used to generate and rearrange an array of laser beams by splitting a single incident laser beam. High-speed optical modulators such as acousto-optic and electro-optic modulators are used for resonant control of the atoms. Combinations of such modulation mechanisms also play an essential role to enlarge the control capabilities. However, development of advanced photonic systems will further extend the range of quantum circuits and simulations implemented on the neutral-atom array systems.
@@
To summarize, quantum control applications require high speed ($>$ MHz) modulation for wavelengths across the UV and visible spectrum. Size, optical loss, and control power will be important for deployable systems and will be important to scale to the millions of channels needed for utility-scale quantum information processing. Although not discussed in detail in this review, optical power handling will also be a particular concern, especially for two-photon control as necessary for Raman single and multi-qubit gates in trapped ions and optical tweezer generation and multi-photon excitation for neutral atoms. Additionally, isolation between on and off states is an important consideration for all of the quantum operations discussed here, with eventual goals of isolation better than 40 dB to reach the gate errors needed for fault-tolerant quantum computation. Both power handling and isolation are figures of merit that are not often optimized for in current integrated visible-light solutions. The importance of the main figures of merit discussed in this review (speed, control power, size, optical bandwidth, and technical maturity) are qualitatively depicted in Fig. 2(a).
@@ I.2 Augmented-Reality Displays
[…]
The field of visible-light integrated photonics has the potential to enable solutions that address this need for next-generation augmented-reality-display technologies. There have been a number of recent proposals and initial passive demonstrations of integrated-photonics-based near-eye displays that utilize holographic image projection to emit full phase fronts and resolve the vergence-accommodation conflict [101, 102, 103, 104]. As these systems progress to demonstrations with dynamic holographic video display functionality, they will require advances in integrated visible-light modulators. Specifically, these integrated-photonics-based augmented-reality displays will require integrated modulators with visible-spectrum-spanning functionality for multi-wavelength RGB operation. As summarized in Fig 2(a), compact size and low-power operation is paramount for densely-integrated, wearable, high-resolution systems, but they will not require high modulation speeds since the refresh rate of the human eye is only approximately 60 Hz.

### v2

# Technologies for Modulation of Visible Light and their Applications

@@ Abstract
Control over the amplitude, phase, and spatial distribution of visible-spectrum light underlies many technologies, but commercial solutions remain bulky, require high control power, and are often too slow. Active integrated photonics for visible light promises a solution, especially with recent materials and fabrication advances. In this review, we discuss three growing application spaces which rely on control of visible light: control and measurement of atomic quantum technologies, augmented-reality displays, and measurement and control of biological systems. We then review the commercial dynamic surfaces and bulk systems which currently provide visible-light modulation and the current state-of-the-art integrated solutions. Throughout the review we focus on speed, control power, size, optical bandwidth, and technological maturity when comparing technologies.
@@
Control over the temporal and spatial characteristics of light has revolutionized fields from communication to the lithography tools which enable the semiconductor industry. In the 19th century, Pockels and others investigated how external electric and strain fields could change the optical properties of materials, enabling a host of bulk optical elements for manipulating light. These bulk modulators are the current workhorse for light modulation with near-unity efficiencies and modulation speeds up to GHz. However, these components provide limited spatial control, are generally large, and require high drive power.
@@
Since the early 1990s, miniaturized and integrated systems have been recognized as the future of active light control [1] and integrated photonics and controllable surfaces promise to reduce both the size and power requirements of visible-light modulation devices. Indeed, for light with frequencies below the bandgap of Silicon (Si) ($<1.12$ eV, $\lambda>1100$ nm) great progress has been made in developing compact, fast modulators based on integrated photonics [2, 3, 4]. The trillion-dollar electronic chip industry relies on complementary metal-oxide semiconductor (CMOS) fabrication, and by leveraging this industry, Si integrated photonics has been able to move from proof-of-principle demonstrations to deployment in data centers and high-rate intercontinental communication [5]. However, the miniaturization and commercialization of devices to control visible light has so far lagged behind.
@@
Control of visible light at low per-channel drive power, compact form-factor, and high frequencies will enable the growth of a wide variety of technologies. In particular, control and measurement of quantum systems, augmented-reality displays, and imaging and control of biological systems are three application areas that have enjoyed significant growth in the last decade, but which are hampered by the capabilities of current commercially-available visible-light modulators. This explosion of application areas coupled with materials and processing developments has accelerated the research in and development of technologies for visible-light modulation. Here, we review the needs of these three application spaces and the current state-of-the-art for visible-light modulation technologies. We aim to provide a review of the application space to motivate and focus further research in scalable visible light modulation platforms and a review of the optical state-of-the-art for researchers in application fields who are interested in scaling up modulation of visible light.
@@
In Sec. I, we discuss the needs of three application spaces in detail - quantum control and readout of atomic qubits, augmented-reality displays, and control of biological systems. These technologies serve as an anchor point for the figures of merit of the various visible-light modulation platforms discussed in the following sections. In Sec. II, we discuss reflective and transmissive controllable surfaces. These, mainly commercial, solutions enable current application systems but suffer from high per-channel power and size, and often cannot reach the necessary modulation speeds. In Sec. III, we discuss integrated solutions to visible-light modulation. These nascent technologies need significant progress to reach the maturity needed for systems-level integration, but hold great promise. Finally, we conclude with a perspective on future research directions needed to transition these technologies out of the laboratory and into the field.
@@ I Applications
Additionally, there are a plethora of other potential applications which may be enabled by improved modulation of visible light. For instance, both classical and quantum optical computing would benefit from moving to visible wavelengths to reduce component size. Similarly, most light detection and ranging (LiDAR) demonstrations use telecommunications light, but visible LiDAR can outperform infrared systems in underwater applications due to reduced absorption [6, 7, 8]. Finally, visible light communication or Li-Fi [9] has applications in indoor, high-accuracy positioning or when electromagnetic interference must be avoided. The wavelengths of interest for all these applications are shown in Fig. 4.
@@ I.1 Quantum Technologies
Quantum technologies promise improvements over their classical counterparts in communication [10], sensing [11], timekeeping [12, 13], and computation [14]. However, to provide utility beyond proof-of-principle demonstrations it will be necessary to reduce the size, weight, and power of all components while simultaneously improving stability and capability. Neutral atom and charged ion qubits are two of the leading platforms for quantum information processing (QIP) and both rely on ultraviolet, visible, and near infrared light for control and measurement. Although we do not discuss them explicitly here, quantum control of molecules [15] and a broad range of solid state atomic-like defects, for example in diamond [16, 17] or SiC [18], also rely on visible light modulation, with requirements similar to neutral atoms and trapped ions.
Figure 1: Block diagram depicting the optical control system for quantum technologies. Laser light is routed through an optical system which splits each source into many channels, modulates each channel, and shapes the light to control the quantum system.
The majority of experimental demonstrations of atomic QIP have relied on bulk crystals to modulate the light needed to control and readout the quantum state of the system. However, to scale these systems to the millions of qubits necessary for fault-tolerant quantum computation [19], it will be necessary to actively modulate the amplitude, frequency, phase, and often position of millions of channels simultaneously at wavelengths across the UV and visible spectrum. This is untenable with current solutions. Similarly, deployable quantum sensors, clocks, and quantum network nodes will need compact, stable optical control. It is thus clear that integrated photonics will play a critical role in future quantum technologies [20, 21, 22].
@@ I.1.1 Trapped Ions
Figure 2: Electronic level structure for 40Ca+, 137Ba+, and 171Yb+ as three representative ions, and for 87Rb and 171Yb as two representative neutral atoms. Each transition is labeled with its function for quantum control. Light is scattered off Readout transitions to measure the quantum state. Cooling transitions are used to reduce the motion of a trapped atomic qubit to near the ground state. Shelving transitions are used to hide some atomic population in along lived state, usually during quantum state discrimination. Repump transitions are used to remove population from a state where it may otherwise be trapped. Rydberg transitions are used to excite an atom to a highly excited state. Tweezer beams are detuned from a dipole transition and are used to confine an atom. Finally, Qubit transitions are used to control the state of the qubit.
Trapped ions are a promising platform for quantum simulation [23, 24], quantum sensing [25, 26, 27], optical clocks [28], and gate-based quantum computation [29, 30, 31]. Indeed, the state-of-the-art systems provide the highest reported quantum volume to date [32, 33]. In trapped-ion quantum technologies, atomic ions (generally singly charged) are trapped in ultrahigh vacuum using electromagnetic fields. Quantum information is stored in the coherence between two stable states of each ion. Each ion qubit is by nature identical and can be well-isolated from the environment. Furthermore, ions trapped together in the same confining potential share motional modes that can be used to directly mediate entanglement between qubits [34], leading to natural all-to-all connectivity within a chain of ions trapped in a single potential. Entanglement between ions in different potentials can be mediated by either physically shuttling the ions between trapping regions [33] or optically [35].
Fig. 2 shows relevant transitions for controlling (a) 40Ca+ and (b) 137Ba+, and (c) 171Yb+, three commonly used ions. Other ions and isotopes exhibit similar level structures, though with different transition frequencies. As evidenced in Fig. 2, control of even a single species can span the spectrum from the ultraviolet to the near infrared. Additionally, it is likely that a final trapped-ion QIP architecture will employ multiple ion species [33] or encodings [36] for mid-circuit cooling and measurement.
Trapped-ion qubits can generally be separated into two categories – optical and ground-state qubits. For example, 40Ca+ optical qubits are encoded between one state in the $4^{2}$1/2 manifold and one state in the metastable $3^{2}$5/2 manifold [37]. This is a weakly-allowed quadrupole transition with an optical transition at 729 nm. Driving single qubit gates requires only modest power ($\sim 10\mu$W per qubit) but stringent frequency and phase stability [29]. Conversely, in 137Ba+ [38] and 171Yb+ [39], qubits are generally encoded in the S1/2 hyperfine levels. These qubits have GHz-scale splittings and can be driven directly with microwave fields or optically through two-photon Raman transitions [40]. Raman control requires high power ($\sim 10$ mW per qubit) and often low wavelengths ($<$ 400 nm), but the requirements for frequency and phase stability are relaxed as only the difference frequency and phase between the two Raman beams must be precisely controlled. For both qubit encodings, single qubit gate times are around 1 $\mu$s, and two qubit gate times are around 100 $\mu$[29], necessitating amplitude, phase, and frequency control at speeds above 1 MHz.
@@
Recently, piezo-actuated waveguide modulators (see Sec. III.4) were used to control the state of a 40Ca+ trapped-ion qubit with MHz switching speeds and nearly 40 dB on-off extinction [60]. With the integrated modulators they measured 99.7 % single qubit gate fidelity with randomized benchmarking and performance that nearly matched the performance of standard acoustic optic modulation. However, they found that the integrated modulators were significantly less stable, highlighting that significant device and systems-level improvements are needed to reach the stringent requirements needed for quantum control.
I.1.2 Neutral Atoms in Optical Tweezer Arrays
@@
Neutral-atom arrays are an another leading platform for quantum metrology [61, 62, 63, 64], simulation [65, 66, 67, 68, 69, 70, 66], and information processing [71, 72, 73, 74, 75]. Recently, they have also enabled impressive advances towards fault-tolerant quantum processing [75]. The neutral atoms are laser-cooled and then loaded into programmable arrays of optical tweezers [76]. Defect-free arrays with hundreds to thousands of atoms are possible [68, 77, 78, 79, 80]. The neutral-atom qubits are two stable internal states of the atoms with a long coherence time. Like ion qubits, they are identical and well-isolated from their environment.
@@
In neutral atom QIP, lasers are used to trap and control the position of individual neutral-atom qubits and mediate single and multi-qubit gates. For entangling gates, laser beams selectively excite target atoms to highly excited Rydberg states, enabling mutual interactions via the Rydberg blockade effect [81, 82]. Nearby atoms cannot be simultaneously promoted to the Rydberg state, naturally implementing the conditional-NOT [83, 72, 71, 82] and Toffoli [84] gates between the qubits. This effect can also be used to control the system’s many-body Hamiltonian for quantum simulation and optimization [85, 67, 68, 73]. Moreover, the dynamic programming of tweezer positions allows the entangled qubits to be transported across an array. This provides reconfigurable, all-to-all connectivity essential for efficiently implementing hard problems [73, 86], quantum error-correction codes [71, 87] and fault-tolerant information processing [88, 89, 75].
@@
Control of neutral-atom qubits uses multiple colors of laser light (Fig. 2). For instance, an array of green laser foci creates dipole trap arrays for 171Yb qubits [90]. Driving the ground-to-Rydberg transition of Rubidium uses 297 nm (the combination of 480 and 780 nm or 420 and 1013 nm) lasers for the single- (two-) photon transition [91, 92, 85]. Laser cooling and trapping neutral atoms requires kHz timing control, and the timescales for quantum gates typically ranges from 1 MHz [65, 70] to 100 MHz [82, 92, 93] depending on the implementation. Moreover, since the atom arrangement determines the qubit connectivity [71, 73, 86] the dynamic control of spatial modes combined with the fast temporal modulation (e.g., controlling the position of the lasers providing the trapping potentials) is essential for assembling and rearranging atom arrays and implementing local qubit control [75].
@@
The required control currently utilizes different modulation mechanisms. For optical tweezers, liquid-crystal-on-silicon (LCoS) spatial light modulators (Section III.5), mechanical mirror arrays (Section II.1), and acousto-optic modulators (Section II.2.1) are used to generate and rearrange an array of laser beams from a single incident laser beam. High-speed optical modulators such as acousto-optic and electro-optic modulators are used for resonant control of the atoms. Combinations of such modulation mechanisms also play an essential role to enlarge the control capabilities. However, development of advanced photonic systems will further extend the range of quantum circuits and simulations implemented on the neutral-atom array systems.
@@
To summarize, quantum control applications require high speed ($>$ MHz) modulation for wavelengths across the UV and visible spectrum. Size, optical loss, and control power will be important for deployable systems and will be important to scale to the millions of channels needed for utility-scale quantum information processing. Although not discussed in detail in this review, optical power handling will also be a particular concern, especially for two-photon control as necessary for Raman single and multi-qubit gates in trapped ions and optical tweezer generation and multi-photon excitation for neutral atoms. Additionally, isolation between on and off states is an important consideration for all of the quantum operations discussed here, with eventual goals of isolation better than 40 dB to reach the gate errors needed for fault-tolerant quantum computation. Both power handling and isolation are figures of merit that are not often optimized for in current integrated visible-light solutions but should beconsidered moving forward. The importance of the main figures of merit discussed in this review (speed, control power, size, optical bandwidth, and technical maturity) are qualitatively depicted in Fig. 4(a).
@@ I.2 Augmented-Reality Displays
Typical commercially-available augmented-reality head-mounted displays employ an optical relay system for each eye, wherein an image produced by a microdisplay is magnified using a system of lenses to generate an image superimposed on the external scene at a single virtual focal plane in the user’s FOV [98, 99]. However, the bulk-optics components utilized in these typical head-mounted displays result in large, heavy, and indiscreet systems. Additionally, head-mounted displays typically employ low-luminance microdisplays (approximately 1000 cd/m2), which render the systems inadequate for use in ambient daylight conditions (though recent developments in high-density microLEDs show promise for improved luminance [100]). Moreover, typical head-mounted displays utilize optical relay systems with limited FOVs (limited to < ${<}\,40^{\circ}$ compared to the 60 near-peripheral FOV of the human eye). Finally, typical head-mounted displays magnify the microdisplay image such that it appears at a single virtual focal plane and are not capable of producing holographic images with full depth cues. This lack of depth information results in users experiencing eyestrain and headaches that limit long-term and wide-spread use of these displays, an effect known as the vergence-accommodation conflict [101, 102, 103]. Given these limitations, there is a growing need for a discrete, mobile, large-FOV, and high-brightness augmented-reality head-mounted display with full binocular and monocular depth cues.
The field of visible-light integrated photonics has the potential to enable solutions that address this need. There have been a number of recent proposals and initial passive demonstrations of integrated-photonics-based near-eye displays that utilize holographic image projection to emit full phase fronts and resolve the vergence-accommodation conflict [104, 105, 106, 107]. As these systems progress to demonstrations with dynamic holographic video display functionality, they will require advances in integrated visible-light modulators. Specifically, as summarized in Fig 4(a), these integrated-photonics-based augmented-reality displays will require integrated modulators with visible-spectrum-spanning functionality for multi-wavelength RGB operation and with compact sizes and low-power operation for densely-integrated high-resolution systems. However, they will not require high modulation speeds since the refresh rate of the human eye is only approximately 60 Hz.